Circular No. 7798

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002H IN MCG –02-35-11 D. Weisz and W. D. Li, University of California at Berkeley, report the

LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Jan. 19.5 (mag ∼ 16.7) and 20.5 UT (mag ∼ 16.8). SN 2002H is located at α = 13^{h}38^{m}27^{s}.58, δ = −10^{o}43′12′′.9 (equinox 2000.0), which is 32′′.0 west and 58′′.0 south of the nucleus of MCG –02-35-11. A KAIT image taken on 2001 June 18.3 showed nothing at this position (limiting mag ∼ 19.0).

X-RAY TRANSIENTS IN M31 S. Trudolyubov, W. Priedhorsky, and K. Borozdin, Los Alamos Na-

tional Laboratory; K. Mason, Mullard Space Science Laboratory; and F. Cordova, University of California, Santa Barbara, report on the discovery of the three new transient x-ray sources in M31: “The transients were detected in 50 000-s exposures with EPIC instruments during XMM-Newton observations on Jan. 5. The first source (XMMU J004414.0+412204), located at α = 0^{h}44^{m}14^{s}.0, δ = +41^{o}22′04′′ (equinox 2000.0; 4′′ error radius), exhibits a supersoft x-ray spectrum (thermal spectrum with kT ∼ 40 eV) with an observed luminosity (0.3–1.5 keV) of 8 × 10^{36} erg/s, assuming a distance of 760 kpc. A second transient source (XMMU J004415.9+413058), is located at α = 0^{h}44^{m}15^{s}.9, δ = +41^{o}30′58′′ (equinox 2000.0; 3′′ error radius), and it exhibits a highly absorbed spectrum [photon index ∼ 1.4 for power-law approximation or kT = 1.4 keV for blackbody approximation and an equivalent hydrogen column density of (0.5–1.4) × 10^{22} cm^{−2}, depending on the continuum model] and observed luminosity (0.2–7 keV) of 8 × 10^{36} erg/s; there is possible evidence for a quasiperiodic modulation of the source x-ray flux with a period of ∼ 5600 s. The third transient source, XMMU J004438.1+414511, is located at α = 0^{h}44^{m}38^{s}.1, δ = +41^{o}45′11′′ (equinox 2000.0; 4′′ error radius); this source was detected at a luminosity of 8 × 10^{36} erg/s (0.2–7 keV), with a spectrum well fitted with a power law of photon index 1.9. The three sources were not detected in earlier Chandra HRC and deep ROSAT observations, down to luminosities 10–20 times lower than measured by XMM. Observations of the three new x-ray transients at other wavelengths are encouraged.”

2002 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green